OpenStudy (anonymous):

What is the electric field on the surface of a thin conducting spherical shell containing uniform charge Q?

OpenStudy (anonymous):

Is it kQ/r^2 or 0?

OpenStudy (anonymous):

@dan815

OpenStudy (rushwr):

\[E=\frac{ 1 }{ 4\pi \epsilon } \frac{ Q }{R ^{2}}\]

OpenStudy (anonymous):

How to derive that?

OpenStudy (irishboy123):

use a Gausssian surface and Gauss's Law [\(\Phi = \int \vec E . \hat n \ da = \frac{ Q_{enc}}{\epsilon_o} \)]. if the outer radius of the shell is R, then \(\large E \times 4 \pi R^2 = \frac{Q}{\epsilon_o}\) -- leading to the formula posted above you can generalise for all radii

OpenStudy (anonymous):

I understand the electric fields inside and outside the shell, but I'm having doubt only about a point on the shell, since it's a thin shell is charge Q enclosed within the Gaussian surface?

OpenStudy (irishboy123):

doesn't really matter if its thin or thick; as long as it's conducting, the charge will be at the surface so looks to me like a practical question.